### v1

@@ Abstract
Semantic-based test generators are widely used to produce failure-inducing inputs for Deep Learning (DL) systems. They typically generate challenging test inputs by applying random perturbations to input semantic concepts until a failure is found or a timeout is reached. However, such randomness may hinder them from efficiently achieving their goal. This paper proposes XMutant, a technique that leverages explainable artificial intelligence (XAI) techniques to generate challenging test inputs. XMutant uses the local explanation of the input to inform the fuzz testing process and effectively guide it toward failures of the DL system under test. We evaluated different configurations of XMutant in triggering failures for different DL systems both for model-level (sentiment analysis, digit recognition) and system-level testing (advanced driving assistance). Our studies showed that XMutant enables more effective and efficient test generation by focusing on the most impactful parts of the input. XMutant generates up to $125\%$ more failure-inducing inputs compared to an existing baseline, up to 7$\times$ faster. We also assessed the validity of these inputs, maintaining a validation rate above $89\%$, according to automated and human validators.
journal: Journal
@@ 1 Introduction
An implementation of our approach for focused fuzzing of DL systems, implemented in the tool XMutant. To the best of our knowledge, this is the first solution that uses XAI techniques for semantic-based focused test generation, both at the model and system level.
@@ 3 XMutant
Testing scenarios.  Table 1 shows the characteristics of our testing scenarios. Semantic representations are often domain-specific. For example, both sentiment analysis and digit recognition are classification tasks, but the semantic modelings are different. In sentiment analysis, the words in a movie review are mapped to an embedding space whereas in digit recognition, the bitmap of a MNIST handwritten digit mnist is converted into sequences of cubic Bézier curves farin1983algorithms defined by a series of control points, according to the Scalable Vector Graphics representation. For advanced driving assistance, the centerline of a road in a driving scenario is represented by Catmull-Rom cubic splines catmull1974class , also specified by a sequence of control points. The different testing levels also imply distinct interpretations of failures. For sentiment analysis and digit recognition, a failure is defined as the disagreement between the output of the DNN under test and the ground truth, e.g., a misclassification of the movie review or the handwritten digit. For advanced driving assistance, failures are identified by the misbehavior of the whole DL system, i.e., the autonomous vehicle, in response to the outputs of the DNN, rather than isolated incorrect predictions on individual images. Following existing research Abdessalem-ICSE18 , we evaluate whether a sequence of incorrect DNN predictions leads to a violation of the safety requirements of the system, i.e., the vehicle driving out of the road, or a reduction of the driving quality 2021-Jahangirova-ICST .
@@ 6             Test DL system with the input $t$;
8                   $S$leftarrow GetSemanticRepresentation($t$) ;
9                   $e$leftarrow LocalExplanationComputation($t$) ;
10                   $cs$leftarrow SemanticConceptSelection($S$, $e$) ;
11                   $d$leftarrow MutationDirectionComputation($S$, $e$, $cs$) ;
12                   $S$leftarrow Mutate($S$, $cs$, $d$) ;
13                   $t$leftarrow GenerateConcreteInputFromSemantic($S$) ;
@@ 3.1 XAI-guided Mutation Operator
XMutant’s mutation operator comprises three steps: (1) computing the local explanation (Line 8), (2) selecting the semantic concept to mutate (Line 9), and (3) determining the mutation direction (Line 10). First, XMutant uses an XAI technique to compute local explanations associated with a DNN and its test instances (procedure LocalExplanationComputation on Line 8). Typically, the explanation of the DNN prediction for a given input has the same size as the input, with each dimension of the input space corresponding to an explanation score. Second, XMutant selects the semantic concept to mutate on the high-attention area of the local explanation (procedure SemanticConceptSelection on Line 9), where an intermediate weight vector ${w}$ is obtained and denotes the importance of semantic concepts with $dim({w})=dim(S)$. For textual inputs, the local explanation corresponds directly to the semantic concept, hence selecting the semantic representation is based on the explanation’s magnitude (i.e., ${w}=e$ and $dim(e)=dim({w})$). However, for model- or system-level imagery inputs, our approach needs to map the high-dimensional explanations, onto the lower-dimensional semantic concept space through a function $f$ before the selection, i.e., ${w}=f(e)$ and $dim(e)>dim({w})$. To ensure diversity of generated test inputs, XMutant randomly selects candidate semantic concepts $cs$ based on the weights ${w}$, where semantic concepts with higher weights have higher chances of being chosen Third, when a candidate semantic concept is selected, XMutant leverages the local explanations to choose the mutation direction, aiming to maximize the mutation effectiveness (Line 10).
XMutant is configured according to the kind of input and semantic representation associated with the task under test (Table 1). We elaborate on the domain-specific configurations in terms of semantic concept selection and mutation direction in the following sections.
@@ 3.2 Model-level Application on Textual Inputs
For textual inputs, the inputs to the DNN are tokens corresponding to the words. Therefore, obtained explanations have the same dimensions as the semantic representation, i.e., words.
With model-agnostic XAI technique methods such as LIME lime , the feature attributions for each token can be computed directly. With gradient-based XAI methods, however, the presence of a non-differentiable embedding layer in the DL model under test, makes the gradient computation unavailable. To address this, we propose a customized approach for interpreting such DL models. Let us assume that the embedding layer mapping the input vector $x$ consisting of $L$ tokens in the embedding space of dimension $N$, is a matrix of size $\mathbb{R}^{L\times N}$. Then we take the following steps. First, we remove the embedding layer from the model under test to obtain the corresponding submodel. Second, we apply gradient-based XAI methods to compute the explanation $\hat{e}\in\mathbb{R}^{L\times N}$ in the embedding space for the submodel, with each row vector $\hat{e}[i]\in\mathbb{R}^{N}$ representing the explanation vector for the corresponding token $x[i]$. Third, we map the explanation in the embedding space back to the input space with summation (for Integrated Gradients) or norm operators (for SmoothGrad), e.g., $e[i]=sum(\hat{e}[i])$ for each token $x[i]$. Finally, we obtain a local explanation $e$ that has the same dimension as the input $x$.
The local explanations obtained for textual inputs differ in their characteristics depending on the XAI method used. For instance, SmoothGrad captures the gradient information, indicating the sensitivity of the predictions to word variations (Figure 1 (a)). In this case, the weight vector consists of absolute gradient information. In contrast, LIME and Integrated Gradients provide explanations on feature attributions, which represent the magnitude of the positive (or negative) contribution of a word on the prediction, making them useful for guiding the mutation direction (Figure 1 (b)). In this case, the weight vector contains the raw explanations.
@@
Regarding the actual mutation operation, we used two mutation methods available in the literature deepatash : (1) replacing a word with its synonym obtained from WordNet miller1995wordnet , and (2) adding an “and” conjunction after an adjective (or adverb), followed by a synonym of the adjective (or adverb). These mutation methods ensure that the original meaning of the sentence is preserved, thereby maintaining a high validity of generated test inputs.
During the fuzzing process, one of the mutation methods is selected at each iteration. When the first mutation method is chosen, we apply it to the selected word, regardless of the considered XAI method. This is because it is uncertain whether replacing a synonym will increase or decrease the contribution of the selected word. In contrast, the second mutation method enhances the semantics of the sentence by adding a synonym. Therefore, for XAI methods that differentiate feature attributions (i.e., LIME lime and Integrated Gradients integratedgradients ), we perform targeted mutation by applying the second mutation method only to words with opposite effects to the prediction (e.g., we modify “promptly” to “promptly and quickly” but we do not modify “overwhelmed” in Figure 1), which helps to more effectively challenge the model under test. For XAI methods that are not focused on explaining the feature attributions (i.e., SmoothGrad), we apply the second mutation method on the selected word, regardless of its effect on the prediction.
@@ 3.3 Model-level Application on Imagery Inputs
For imagery inputs, the local explanation, so-called heatmap, is itself an image with the same dimensions as the original image, where pixels are highlighted based on the attention that the DNN under test gives to each portion of the image. As such, both the selection of the candidate semantic concept to be mutated and the computation of the mutation direction are based on a single heatmap $\mathbf{H}$.
@@
Selecting the candidate semantic concept, i.e., the control point $cp_{k}$, for mutation requires the computation of the weights corresponding to the control points based on the heatmap. We evaluate two techniques in this paper, based on square windows and clustering.
@@ 3.3.1 Square Windows Semantic Concept Selection
$w_{k}=\frac{1}{(2d+1)^{2}}\sum_{i=-d}^{i=d}\sum_{j=-d}^{i=d}\mathbf{H}[(\mathbf{p}_{c_{k}}+(i,j))]$ (1)
where $d$ is the distance between the center of the square $(x_{k},y_{k})$ and each side (and the window size ws is given by $\textit{ws}=2d+1$), while $\mathbf{p}_{c_{k}}\in\mathbb{R}^{2}$ denotes the coordinates of the control point in the heatmap. Subsequently, the weights ${w}$ for each semantic concept are normalized.
The square windows approach, although straightforward, involves choosing an appropriate window size ws. When the window size is large, it can exceed the input’s boundaries or have squares of different control points overlap. If the window size is small, large regions of the local explanation might be uncovered, potentially missing some high-attention areas. Additionally, this method does not provide any guidance for determining the direction of mutation.
@@ 3.3.2 Clustering-based Semantic Concept Selection
To overcome the limitations of the square windows approach, we also propose a control point selection technique based on clustering, which is less reliant on hyperparameters and, thus, more automated.
XMutant uses the k-means clustering algorithm wu2012advances to cluster pixels based on their values with the number of clusters $n$ to match the number of control points. We modified the typical k-means process by initializing the centroids with the control points and limited the algorithm to a single iteration, without convergence. Figure 2 (d) shows 13 clusters identified by different colors, one for each control point. Then, XMutant computes the weight $w_{k}$ for each control point $cp_{k}$ as the sum of the ratios between each pixel value in the cluster of $cp_{k}$, and the distance between the coordinates of the pixel and the respective control point:
$w_{k}=\sum_{i=1}^{m_{k}}\frac{\mathbf{H}[\mathbf{p}_{i}^{k}]}{d_{i}^{k}},\text{ with }d_{i}^{k}=\max(1,||\mathbf{p}_{i}^{k}-\mathbf{p}_{cp_{k}}||)$ (2)
@@
Once the weights have been normalized and the control point $cp_{k}$ sampled (e.g., the control pointing the green cluster in the bottom-left corner of Figure 2 (e)), XMutant mutates the selected control point towards the attention attractor $\mathbf{c}_{k}$, which is defined as the center of intensity of the respective cluster, given by:
$\mathbf{c}_{k}=\frac{\sum_{i=1}^{m_{k}}\mathbf{H}[\mathbf{p}^{k}_{i}]\mathbf{p}_{i}^{k}}{\sum_{i=1}^{m_{k}}\mathbf{H}[\mathbf{p}^{k}_{i}]}$ (3)
@@ 3.4 System-level Application on Logical ADAS Scenarios
As the ADAS processes a stream of images from a driving simulator, applying our approach at an individual image level (as described previously for model-level testing) does not effectively assist in selecting and mutating candidate semantic concepts, i.e., control points. Instead, XMutant generates a series of heatmaps for each image perceived by the DNN, and aggregates them to retrieve a semantic score that reflects the system’s overall performance.
@@
Indeed, existing work 2022-Stocco-ASE has shown that the derivative of a local explanation is indicative of a poor-performing DNN. In this way, XMutant grants a higher weight to road sections—and associated control points—where the driving behavior of the DNN is more erratic. Consequently, by mutating such control points in the corresponding road sections, XMutant is more likely to generate a test input that induces more challenging driving conditions. For instance, in Figure 3 (d), the road section with the highest mean derivative is road section III; in this example, the corresponding control point in the top-right corner of Figure 3 (e) is sampled.
Ultimately, XMutant determines the mutation direction by considering the perpendicular line to the road’s centerline at the candidate control point. Consequently, only two directions are possible: moving the control point toward the right lane or the left lane.
To determine the appropriate mutation direction, we analyze the series of heatmaps for the chosen road segment. XMutant partitions each local explanation into two equally-sized sub-heatmaps, i.e., $\mathbf{H}[:w/2,:]$ and $\mathbf{H}[w/2+1:,:]$ (where the column separates start and end indices for each dimension of the local explanation). XMutant then calculates the difference in average intensity between the left and right sub-heatmaps in the road segment. The mutation direction is chosen based on the side with the highest average intensity (in Figure 3 (e) the selected direction is the right direction, hence the last control point is translated towards the right lane, resulting in the test input shown in Figure 3 (f)).

### v2

@@ Abstract
[…]
journal: Empirical Software Engineering
@@ 1 Introduction
An implementation of our approach for focused fuzzing of DL systems, implemented in the publicly available tool XMutant replication-package. To the best of our knowledge, this is the first solution that uses XAI techniques for semantic-based focused test generation, both at the model and system level.
@@ 3 XMutant
Testing scenarios.  Table 1 shows the characteristics of our testing scenarios. Semantic representations are often domain-specific. For example, both sentiment analysis and digit recognition are classification tasks, but the semantic modelings are different. In sentiment analysis, the words in a movie review are mapped to an embedding space whereas in digit recognition, the bitmap of a MNIST handwritten digit mnist is converted into sequences of cubic Bézier curves farin1983algorithms defined by a series of control points, according to the Scalable Vector Graphics representation. For advanced driving assistance, the centerline of a road in a driving scenario is represented by Catmull-Rom cubic splines catmull1974class, also specified by a sequence of control points.
The different testing levels also imply distinct interpretations of failures. For sentiment analysis and digit recognition, a failure is defined as the disagreement between the output of the DNN under test and the ground truth, e.g., a misclassification of the movie review or the handwritten digit. For advanced driving assistance, failures are identified by the misbehavior of the whole DL system, i.e., the autonomous vehicle, in response to the outputs of the DNN, rather than isolated incorrect predictions on individual images. Following existing research Abdessalem-ICSE18, we evaluate whether a sequence of incorrect DNN predictions leads to a violation of the safety requirements of the system, i.e., the vehicle driving out of the road, or a reduction of the driving quality 2021-Jahangirova-ICST.
Notably, unlike evolutionary fuzzing approaches, XMutant does not employ a continuous fitness score to rank or retain test inputs. Instead, it solely relies on a binary failure criterion to decide whether to keep a mutated input. The intention is to isolate and evaluate the impact of XAI-guided mutation alone, without the confounding effects of fitness-based selection. This design also preserves flexibility for integrating XMutant into larger testing frameworks that may employ their own prioritization or fitness metrics.
@@ 6     Test DL system with the input $t$;
8        $S$ $\leftarrow$ GetSemanticRepresentation($t$) ;
9        $e$ $\leftarrow$ LocalExplanationComputation($t$) ;
10        cscs $\leftarrow$ SemanticConceptSelection($S$, $e$) ;
11        $d$ $\leftarrow$ MutationDirectionComputation($S$, $e$, cscs) ;
12        $S$ $\leftarrow$ Mutate($S$, cscs, $d$) ;
13        $t$ $\leftarrow$ GenerateConcreteInputFromSemantic($S$) ;
@@ 3.1 XAI-guided Mutation Operator
XMutant’s mutation operator comprises three steps: (1) computing the local explanation (Line 8), (2) selecting the semantic concept to mutate (Line 9), and (3) determining the mutation direction and applying mutation (Line 10, 11). First, XMutant uses an XAI technique to compute local explanations associated with a DNN and its test instances (procedure LocalExplanationComputation on Line 8). Typically, the explanation of the DNN prediction for a given input has the same size as the input, with each dimension of the input space corresponding to an explanation score. Second, XMutant selects the semantic concept to mutate on the high-attention area of the local explanation (procedure SemanticConceptSelection on Line 9), where an intermediate weight vector w{w} is obtained and denotes the importance of semantic concepts with dim(w)=dim(S)dim({w})=dim(S). For textual inputs, the local explanation corresponds directly to the semantic concept, hence selecting the semantic representation is based on the explanation’s magnitude (i.e., w=e{w}=e and dim(e)=dim(w)dim(e)=dim({w})). However, for model- or system-level imagery inputs, our approach needs to map the high-dimensional explanations, onto the lower-dimensional semantic concept space through a function $f$ before the selection, i.e., w=f(e){w}=f(e) and dim(e)>dim(w)dim(e)>dim({w}). To ensure diversity of generated test inputs, XMutant randomly selects candidate semantic concepts $cs$ based on the weights w{w}, where semantic concepts with higher weights have higher chances of being chosen. Third, when a candidate semantic concept is selected, XMutant leverages the local explanations to choose the mutation direction, aiming to maximize the mutation effectiveness (Line 10).
XMutant applies the same XAI-guided fuzzing workflow across different DL applications, and it is configured according to the kind of input and semantic representation associated with the task under test (Table 1). We elaborate on the domain-specific configurations in terms of semantic concept selection and mutation direction in the following sections.
@@ 3.2 Model-level Application on Textual Inputs
For textual inputs, the inputs to the DNN are tokens corresponding to the words. Therefore, obtained explanations have the same dimensions as the semantic representation, i.e., words (Algorithm 1: GetSemanticRepresentation).
With model-agnostic XAI technique methods such as LIME lime, the feature attributions for each token can be computed directly. With gradient-based XAI methods, however, the presence of a non-differentiable embedding layer in the DL model under test, makes the gradient computation unavailable. To address this, we propose a customized approach for interpreting such DL models. Let us assume that the embedding layer mapping the input vector $x$ consisting of $L$ tokens in the embedding space of dimension $N$, is a matrix of size $\mathbb{R}^{L\times N}$. Then we take the following steps. First, we remove the embedding layer from the model under test to obtain the corresponding submodel. Second, we apply gradient-based XAI methods to compute the explanation $\hat{e}\in\mathbb{R}^{L\times N}$ in the embedding space for the submodel, with each row vector $\hat{e}[i]\in\mathbb{R}^{N}$ representing the explanation vector for the corresponding token $x[i]$. Third, we map the explanation in the embedding space back to the input space with summation (for Integrated Gradients) or norm operators (for SmoothGrad), e.g., $e[i]=sum(\hat{e}[i])$ for each token $x[i]$. This process yields the local explanation $e$ with the same dimension as the input $x$, which XMutant later uses to guide mutations (Algorithm 1: LocalExplanationComputation).
The explanations are used to derive a weight vector $w$ over tokens, from which XMutant samples one candidate concept $cs$ (Algorithm 1: SemanticConceptSelection). However, the local explanations obtained for textual inputs differ in their characteristics depending on the XAI method used. For instance, SmoothGrad captures the gradient information, indicating the sensitivity of the predictions to word variations (Figure 1 (a)). In this case, the weight vector consists of absolute gradient information. In contrast, LIME and Integrated Gradients provide explanations on feature attributions, which represent the magnitude of the positive (or negative) contribution of a word on the prediction, making them useful for guiding the mutation direction (Figure 1 (b)). In this case, the weight vector contains the raw explanations.
@@
Regarding the actual mutation operation, we used two mutation methods available in the literature deepatash: (1) replacing a word with its synonym obtained from WordNet miller1995wordnet, and (2) adding an “and” conjunction after an adjective (or adverb), followed by a synonym of the adjective (or adverb). These operators are used to perturb the selected token $cs$, resulting in a modified semantic input $S$ (Algorithm 1: Mutate). These mutation methods ensure that the original meaning of the sentence is preserved, thereby maintaining a high validity of generated test inputs.
During the fuzzing process, one of the mutation methods is selected at each iteration. When the first mutation method is chosen, we apply it to the selected word $cs$, regardless of the considered XAI method. This is because it is uncertain whether replacing a synonym will increase or decrease the contribution of the selected word. In contrast, the second mutation method enhances the semantics of the sentence by adding a synonym. Depending on the explanation, XMutant decides whether to apply directional mutation (e.g., amplifying or attenuating polarity) or random variation (Algorithm 1: MutationDirectionComputation). Therefore, for XAI methods that differentiate feature attributions (i.e., LIME lime and Integrated Gradients integratedgradients), we perform targeted mutation by applying the second mutation method only to the word with opposite effects to the prediction (e.g., we modify “promptly” to “promptly and quickly” but we do not modify “overwhelmed” in Figure 1), which helps to more effectively challenge the model under test. For XAI methods that are not focused on explaining the feature attributions (i.e., SmoothGrad), we apply the second mutation method on the selected word, regardless of its effect on the prediction.
After mutation, the modified semantic input is projected back into the concrete input space—i.e., transformed into a full sentence—which serves as a new test case for the DL system (Algorithm 1: GenerateConcreteInputFromSemantic).
@@ 3.3 Model-level Application on Imagery Inputs
For imagery inputs, the semantic representation $S$ is extracted by converting the bitmap image into a set of SVG control points (Algorithm 1: GetSemanticRepresentation). The local explanation, so-called heatmap, is itself an image with the same dimensions as the original image, where pixels are highlighted based on the attention that the DNN under test gives to each portion of the image (Algorithm 1: LocalExplanationComputation). The subsequent selection of the candidate semantic concept to be mutated and the computation of the mutation direction are based on a single heatmap $\mathbf{H}$.
@@
Selecting the candidate semantic concept $cs$, i.e., the control point $cp_{k}$, for mutation requires the computation of the weights corresponding to the control points based on the heatmap (Algorithm 1: SemanticConceptSelection). We evaluate two techniques in this paper, based on square windows and clustering.
@@ 3.3.1 Square Windows Semantic Concept Selection
$w_{k}=\frac{1}{(2d+1)^{2}}\sum_{i=-d}^{i=d}\sum_{j=-d}^{i=d}\mathbf{H}[(\mathbf{p}_{c_{k}}+(i,j))],$ (1)
where $d$ is the distance between the center of the square $(x_{k},y_{k})$ and each side (and the window size ws is given by $\textit{ws}=2d+1$), while $\mathbf{p}_{c_{k}}\in\mathbb{R}^{2}$ denotes the coordinates of the control point in the heatmap. Subsequently, the weights w{w} for each semantic concept are normalized.
The square windows approach, although straightforward, involves choosing an appropriate window size ws. When the window size is large, it can exceed the input’s boundaries or have squares of different control points overlap. If the window size is small, large regions of the local explanation might be uncovered, potentially missing some high-attention areas. Additionally, this method does not provide any guidance for the direction of mutation, thus the direction is determined randomly (Algorithm 1: MutationDirectionComputation).
@@ 3.3.2 Clustering-based Semantic Concept Selection
To reduce reliance on hyperparameters and offer better spatial coherence, we also propose a control point selection technique based on clustering. XMutant uses the k-means clustering algorithm wu2012advances to cluster pixels based on their values with the number of clusters $n$ to match the number of control points. We modified the typical k-means process by initializing the centroids with the control points and limited the algorithm to a single iteration, without convergence. Figure 2 (d) shows 13 clusters identified by different colors, one for each control point. Then, XMutant computes the weight $w_{k}$ for each control point $cp_{k}$ as the sum of the ratios between each pixel value in the cluster of $cp_{k}$, and the distance between the coordinates of the pixel and the respective control point:
[…]
@@
Once the weights have been normalized and the control point $cp_{k}$ sampled (e.g., the control pointing the green cluster in the bottom-left corner of Figure 2 (e)), XMutant mutates the selected control point towards the attention attractor $\mathbf{c}_{k}$ (Algorithm 1: MutationDirectionComputation), which is defined as the center of intensity of the respective cluster, given by:
$\mathbf{c}_{k}=\frac{\sum_{i=1}^{m_{k}}\mathbf{H}[\mathbf{p}^{k}_{i}]\mathbf{p}_{i}^{k}}{\sum_{i=1}^{m_{k}}\mathbf{H}[\mathbf{p}^{k}_{i}]}.$ (3)
Finally, the new input is obtained by mutating $cs$ toward the attractor $\mathbf{c}_{k}$ (Algorithm 1: Mutate, GenerateConcreteInputFromSemantic).
@@ 3.4 System-level Application on Logical ADAS Scenarios
As the ADAS processes a stream of images from a driving simulator, applying our approach at an individual image level (as described previously for model-level testing) does not effectively assist in selecting and mutating candidate semantic concepts, i.e., control points. Instead, XMutant generates a series of heatmaps for each image perceived by the DNN, and aggregates them to retrieve a semantic score that reflects the system’s overall performance. This allows the selection of critical semantic components for mutation, following the same fuzzing loop described in Algorithm 1.
@@
Indeed, existing work 2022-Stocco-ASE has shown that the derivative of a local explanation is indicative of a poor-performing DNN. In this way, XMutant grants a higher weight to road sections—and associated control points—where the driving behavior of the DNN is more erratic (Algorithm 1: SemanticConceptSelection). Consequently, by mutating such control points in the corresponding road sections, XMutant is more likely to generate a test input that induces more challenging driving conditions. For instance, in Figure 3 (d), the road section with the highest mean derivative is road section III; in this example, the corresponding control point in the top-right corner of Figure 3 (e) is sampled.
Next, XMutant determines the mutation direction by considering the perpendicular line to the road’s centerline at the candidate control point. Consequently, only two directions are possible: moving the control point toward the right lane or the left lane (Algorithm 1: MutationDirectionComputation). To determine the appropriate mutation direction, we analyze the series of heatmaps for the chosen road segment. XMutant partitions each local explanation into two equally-sized sub-heatmaps, i.e., $\mathbf{H}[:w/2,:]$ and $\mathbf{H}[w/2+1:,:]$ (where the column separates start and end indices for each dimension of the local explanation). XMutant then calculates the difference in average intensity between the left and right sub-heatmaps in the road segment. The mutation direction is chosen based on the side with the highest average intensity (in Figure 3 (e) the selected direction is the right direction). The selected control point is then shifted laterally toward that direction (Algorithm 1: Mutate), resulting in a new and more challenging road configuration (Algorithm 1: GenerateConcreteInputFromSemantic), as visualized in Figure 3 (f).